# A Case Study on Misinformation Detection

### v3

@@ 5.2 Detection of suspicious accounts
We reproduce results from a study published in 2022 by Saeed et al. [361]. In summary, the authors propose a method, called TrollMagnifier, for the identification of Reddit accounts that exhibit troll-like behaviors. Like all other account-scoped methods we analyze in the security literature, TrollMagnifier is trained on posting and reply statistics for non-troll and known Russian troll accounts (as identified by Reddit) [361].
Reproducibility of published results. Study authors provided us with pre-processed datasets and classifier code upon email request. The full Reddit Pushshift dataset is freely available online [421]. We were able to replicate original paper results using these materials. Original account handles were anonymized; as such, we were unable to manually verify if the accounts identified in the original study appeared to be troll-like.
@@
Tautological targets. As described in preceding sections (Section 4.3), suspicious account detection suffers from a lack of clear and consistent definitions. Troll accounts cannot be described by degree of automation (while some trolls are bot-like, many others are operated by humans) [367] or the nature of the information they spread (this might be political misinformation, ads, or anything in between) [422]; as such, the clearest definition that Saeed et al. implicitly offer is that a troll is an account that exhibits troll-like behavior: i.e., interacts with known troll accounts, or appears on the same posts or message threads as these accounts. The authors note in their own work that this approach cannot be used to detect novel trolling behaviors, and requires a seed list of known troll accounts for every new detection task (we note in our discussion of Account-scoped data curation practices that the process of identifying these seed accounts is actually a fairly manual one).
Implied versus actual data dimensionality. While the original TrollMagnifier paper strongly implied that the proposed method would leverage networked behaviors to identify troll accounts acting in coordination online, we found that, in actuality, the features under analysis lacked any sort of temporal component and were limited in scope. There were six features in all (described in full in Figure 2), each corresponding to an aggregate engagement statistic. Longitudinal data and timestamps were not available; as such, it was not possible to perform time series analysis. Additionally, account names were not available, which disallowed construction of user graphs.
Feature importance. In our partial dependence analysis, we find that feature x1—commenting statistics—was the sole feature that consistently produced classification accuracy greater than 0.6 (most other features had accuracy no better than random; two features, x2 and x6, had performance below 0.5). Per our earlier observation that many account-scoped methods target behaviors that are difficult to distinguish from routine online activity, we recommend that feature engineering for account- and network-scoped methods reflect some intuition about the nature of actually suspicious behaviors. Furthermore, the performance of the current feature set—no better than random—suggests that manual classification might be as effective as (or even more effective than) an automated approach that only detects a content-agnostic heuristic.
@@ 5.3 Detection of suspicious websites
We reproduce results from a study published in 2018 by Baly et al [43]. In summary: the authors propose a multimodal approach to the classification of news websites. This method is particularly representative of works within this information scope: 50% of works within this scope use a similar mixed-modalities approach to detecting misinformation websites, and Baly et al. include site-specific feature types, including domain and traffic-based features, in their analysis. Baly et al. analyzed website contents, associated social media accounts, and Wikipedia pages in order to perform two classification tasks: fact and political bias classification. The authors developed a dataset of 1066 websites manually labeled for their political leaning (extreme-left, left, left-center, center, right-center, right, and extreme-right) and degree of credibility (low, mixed, high). These labels were extracted from the Media Bias/Fact Check (MBFC) database [311]. We were able to reproduce original study results and perform ablation analyses on existing datasets. We were unable to run the method on an updated dataset, as feature extraction code was not available.
@@
Multimodal features: help or hindrance? We performed an ablation study of the method on the EMNLP18 dataset and analyzed the method’s bias and fact-checking classification functions separately [43]. Specifically, we stratified the original EMNLP18 dataset by political leaning and credibility, as labeled by MBFC, and analyzed the performance of 1) the full feature set, 2) individual features and 3) ablated feature sets (removing one feature type per test). Our results are summarized in Section A. We find that, on 11 out of 12 test datasets, classifier performance using only text-based features (articles and wikipedia, derived from articles randomly sampled from the website in question, and the site’s corresponding Wikipedia page, respectively) was comparable to performance on the full feature set. On five out six datasets, bias classification accuracy on text-only features actually outperformed bias classification on the whole feature set (see the bottom half of Section A), suggesting that the full-site classifier of Baly et al. was effectively a text content classifier.
@@ 6 DISCUSSION
RQ1: Fit. Very few methods that claimed to detect misinformation performed actual fact verification: Instead, they targeted proxy signals that were frequently steps away from promised detection targets. These differences were particularly noticeable in methods that relied heavily on text- and network-based features to perform classification. In those cases, semantic/syntactic signatures and propagation patterns served as proxies for the existence of misinformation. We demonstrated, through our own replication studies, that it is easy to circumvent approaches that rely on style-based cues to perform proxy detection. Additionally, in the absence of current or complete social media post data, we argued that most network-based methods are poorly-equipped to perform early detection of false rumors—this is by design, as the propagation patterns these methods purport to detect are frequently only noticeable after virality.
@@
RQ2: Data curation and model explainability. Lack of access to current, well-annotated datasets remains a serious problem for current and future misinformation research. Across existing datasets, taxonomies for classifying misinformation were inconsistent. Testing on contemporaneous data was uncommon among those papers we annotated, and testing in real-time settings was even rarer. Proof-of-concept experiments oftentimes did not control for data dependencies. Authors describing black-box methods—particularly those employing neural nets or other forms of unsupervised learning—did not disclose feature sets retrieved by their methods. Model explainability for these approaches—if available at all—was very limited. For methods employing multimodal feature sets, temporal consistency across disparate feature sets was poorly controlled, and feature normalization was often questionable.
@@
RQ3: Reproducibility. We noted widespread code and data availability issues: Fewer than 30% of our attempts to locate code and data, or obtain this information from authors, were successful. The code and datasets we were able to retrieve were frequently unusable or out of date. In fact, we were able to reproduce and replicate results on published and current data for only one of our replication studies. In that single case, we found that the article-scoped method performed no better than random (0.50 accuracy) on a current, mixed-domain dataset (Section 5.1). Additionally, for the significant number of methods that purported to detect false rumors on Twitter, the shutdown of the Twitter Academic API poses an existential risk to reproducibility of results and usefulness of dehydrated datasets.
@@
RQ4: Generalizability. Methods that were trained on single-source or single-domain datasets appeared to perform well on data from the same source, or within the same domain; these methods, unsurprisingly, performed poorly on multi-source or out-of-domain topics. These discrepancies are closely tied to undisclosed or uncontrolled-for data dependencies, which we discuss in RQ2. Though we did not attempt to test detection methods on out-of-scope data—e.g., we did not test claim-scoped methods on datasets of whole news articles—it seems unlikely that these methods would fare well in out-of-scope contexts, particularly in light of their already-poor performance on in-scope datasets.
6.1 Future Directions for Research
Following our replication analyses and survey of the academic literature, we offer the following suggestions for possible future directions for research.
What works? Though the majority of this project has been devoted to critique of an existing body of work, we choose to highlight here a number of approaches that we found especially promising. In general, we found methods that 1) clearly described their detection targets and heuristics and 2) provided context for how their detection outcomes might be employed in service of fact-checking work the most methodologically sound [108, 17, 130]. As we note in our survey of commercial fact-checking services, the preponderance of high-volume checking work is hybrid, with (e.g.) potentially problematic content surfaced by automated means and ultimately vetted by human moderators. We recommend that future research efforts be devoted to understanding how best to facilitate collaboration between human and machine moderation techniques—in particular, how automated methods can support human moderators.
Definitive, rather than descriptive, targets. We discuss the knock-on effects of a basic lack of agreement about what does/n’t constitute misinformation throughout the main body of this work: In the absence of a well-defined detection target, detection methods detect proxy signals instead. Though the task of defining misinformation should likely not be left solely to computer scientists, we can nonetheless set our sights on more manageable targets, such as the three below.
@@ 6.1 Future Directions for Research
Nuanced detection tasks. We note that, in general, the detection methods we consider in this work elide subtleties that are particular to the medium under consideration: for instance, article-level detection methods generally apply broad ‘true’/‘false’ classifications to a text under analysis where only single sentences or turns of phrase might be slightly inaccurate. Additionally, the language signals that most methods expressly detect are fairly unsubtle: Suggestion, insinuation, and leading questions are powerful rhetorical tools that might render a newsreader more susceptible to actual misinformation, or that might suggest misinformative ideas via indirect means; no works within our literature corpus expressly targeted these forms of language, however.
Multimedia detection. The detection methods we consider for this work are solely text-based; we note, however, that a good deal of social media misinformation is transmitted via audio, image, and video. This problem has become especially salient in the current year, as election misinformation transmitted via audiovisual means is rampant on messaging apps worldwide.
Designing for evasion resistance. The AI/NLP literature we review aligns with the current dogma in those subfields, which emphasizes generalizability of single models to all detection settings. We argue that a security framework for method development—wherein attempts at evasion and attacks are assumed, and iterative development is required to stay a step ahead of adversaries—is a more practical approach to misinformation detection research, and will yield more durable outcomes in the long run.

### v4

@@ 5.2 Detection of suspicious accounts
We reproduce results from a study published in 2022 by Saeed et al. [saeed2022trollmagnifier]. In summary, the authors propose a method, called TrollMagnifier, for the identification of Reddit accounts that exhibit troll-like behaviors. Like all other account-scoped methods we analyze in the security literature, TrollMagnifier is trained on posting and reply statistics for non-troll and known Russian troll accounts (as identified by Reddit) [saeed2022trollmagnifier].
Reproducibility of published results. Study authors provided us with pre-processed datasets and classifier code upon email request. The full Reddit Pushshift dataset is freely available online [baumgartner2020pushshift]. We were able to replicate original paper results using these materials. Original account handles were anonymized; as such, we were unable to verify if the accounts identified in the original study appeared to be troll-like.
@@
Tautological targets. As described in preceding sections (Section 4.3), suspicious account detection suffers from a lack of clear and consistent definitions. Troll accounts cannot be described by degree of automation (while some trolls are bot-like, many others are operated by humans) [cresci_bots_2020] or the nature of the information they spread (this might be political misinformation, ads, or anything in between) [yang2022botometer]; as such, the clearest definition that Saeed et al. implicitly offer is that a troll is an account that exhibits troll-like behavior: i.e., interacts with known troll accounts, or appears on the same posts or message threads as these accounts. The authors note in their own work that this approach cannot be used to detect novel trolling behaviors, and requires a seed list of known troll accounts for every new detection task.
Implied versus actual data dimensionality. While the original TrollMagnifier paper strongly implied that the proposed method would leverage networked behaviors to identify troll accounts acting in coordination online, we found that, in actuality, the features under analysis lacked any sort of temporal component and were limited in scope. There were six features in all (described in full in Figure 2), each corresponding to an aggregate engagement statistic. Longitudinal data and timestamps were not available; as such, it was not possible to perform time series analysis. Account names were not available, disallowing construction of user graphs.
Feature importance. In our partial dependence analysis, we find that feature x1—commenting statistics—was the sole feature that consistently produced classification accuracy greater than 0.6 (most other features had accuracy no better than random). Per our earlier observation that many account-scoped methods target behaviors that are difficult to distinguish from routine online activity, we recommend that feature engineering for account- and network-scoped methods reflect some intuition about the nature of actually suspicious behaviors. Furthermore, the performance of the current feature set suggests that manual classification might be as effective as (or even more effective than) an automated approach that detects a content-agnostic heuristic.
@@ 5.3 Detection of suspicious websites
We reproduce results from a study published in 2018 by Baly et al [baly_predicting_nodate]. In summary: the authors propose a multimodal approach to the classification of news websites. This method is particularly representative of works within this information scope: 50% of works within this scope use a similar mixed-modalities approach to detecting misinformation websites, and Baly et al. include site-specific feature types, including domain and traffic-based features, in their analysis. Baly et al. analyzed website contents, associated social media accounts, and Wikipedia pages in order to perform two classification tasks: fact and political bias classification. The authors developed a dataset of 1066 websites manually labeled for their political leaning (extreme-left, left, left-center, center, right-center, right, and extreme-right) and degree of credibility (low, mixed, high). These labels were extracted from the Media Bias/Fact Check (MBFC) database [mediabiasfactheck]. We were able to reproduce original study results and perform ablation analyses on existing datasets. We were unable to run the method on an updated dataset, as feature extraction code was not available.
@@
Multimodal features: help or hindrance? We performed an ablation study of the method on the EMNLP18 dataset and analyzed the method’s bias and fact-checking classification functions separately [baly_predicting_nodate]. Specifically, we stratified the original EMNLP18 dataset by political leaning and credibility, as labeled by MBFC, and analyzed the performance of 1) the full feature set, 2) individual features and 3) ablated feature sets (removing one feature type per test). Our results are summarized in Appendix A. We find that, on 11 out of 12 test datasets, classifier performance using only text-based features (articles and wikipedia, derived from articles randomly sampled from the website in question, and the site’s corresponding Wikipedia page, respectively) was comparable to performance on the full feature set. On five out six datasets, bias classification accuracy on text-only features actually outperformed bias classification on the whole feature set (see the bottom half of Appendix A), suggesting that the full-site classifier of Baly et al. was effectively a text content classifier.
@@ 6 DISCUSSION
RQ1: Fit. Very few methods that claimed to detect misinformation performed actual fact verification: Instead, they targeted proxy signals that were frequently steps away from promised detection targets. These differences were particularly noticeable in methods that relied heavily on text- and network-based features to perform classification. In those cases, semantic/syntactic signatures and propagation patterns served as proxies for the existence of misinformation. We demonstrated, through our own replication studies, that it is easy to circumvent approaches that rely on style-based cues to perform proxy detection.
@@
RQ2: Data curation and model explainability. Lack of access to current, well-annotated datasets remains a serious problem for current and future misinformation research. Across existing datasets, taxonomies for classifying misinformation were inconsistent. Testing on contemporaneous data was uncommon among those papers we annotated, and testing in real-time settings was even rarer. Proof-of-concept experiments oftentimes did not control for data dependencies. Authors describing black-box methods—particularly those employing neural nets or other forms of unsupervised learning—did not disclose feature sets retrieved by their methods.
@@
RQ3: Reproducibility. We noted widespread code and data availability issues: Fewer than 30% of our attempts to locate code and data, or obtain this information from authors, were successful. The code and datasets we were able to retrieve were frequently unusable or out of date. In fact, we were able to reproduce and replicate results on published and current data for only one of our replication studies. In that single case, we found that the article-scoped method performed no better than random (0.50 accuracy) on a current, mixed-domain dataset (Section 5.1).
@@
RQ4: Generalizability. Methods that were trained on single-source or single-domain datasets appeared to perform well on data from the same source, or within the same domain; these methods, unsurprisingly, performed poorly on multi-source or out-of-domain topics. These discrepancies are closely tied to undisclosed or uncontrolled-for data dependencies, which we discuss in RQ2.
6.1 Recommendations for future research
We propose the following as directions for future research:
Designing for intervention fit. All article-scoped methods detect proxy targets for misinformation. Some of these proxy targets, including house style and text sentiment, are easily circumvented (Section 5.1). At the point of task formulation, researchers might consider the following: What signals—or proxy signals—will the intervention detect? What are the potential risks of a false negative or false positive result (and are they worth the potential benefits of e.g., greater speed)? What is the actual detection target, and how closely does it approximate the classification task to be performed? In an adversarial setting, would a malicious actor be able to circumvent the automation?
Understanding hybrid detection. While academic misinformation detection favors fully automated methods, commercial checking services and online platforms employ hybrid methods. Increasingly, online services are turning toward human-driven hybrid detection approaches, wherein human users make complex fact-checking decisions and automated methods amplify their decision-making power (e.g., by identifying content similar to already-flagged posts). We believe that exploration of the interplay between human and automated decision-making systems will prove a fruitful frontier for academic research. And, in fact, recent work [sehat2024misinformation, lykouris2024learning] frames the hybrid detection problem as one of triage and scheduling: how can automated approaches quickly surface urgent cases for further review by human moderators?
Investigating distributional shifts. Fewer than 10% of methods within our whole corpus tested on contemporaneous data, and fewer than 5% tested in a real-time setting. To understand how an ML intervention might work in practice, researchers must understand how robust their methods are to changes in the distribution of available data during training versus testing. These shifts are particular to medium (for instance, whole site classifiers are susceptible to fluctuations in news vertical coverage and author). Notably, the few works in our corpus that tested in real-time reported precipitous performance dropoffs.
@@ 6.1 Recommendations for future research
Addressing nuanced detection challenges. We note that, in general, the detection methods we consider in this work elide subtleties that are particular to the medium under consideration: for instance, article-level detection methods generally apply broad ‘true’/‘false’ classifications to a text under analysis where only single sentences might be slightly inaccurate. Additionally, the language signals that most methods expressly detect are fairly unsubtle, and are likely verifiable via manual means. Suggestion, insinuation, and leading questions are powerful rhetorical tools that might render a newsreader more susceptible to actual misinformation, or that might suggest misinformative ideas via indirect means; no works within our literature corpus expressly targeted these forms of language, however.
7 Ethics considerations
Our author outreach survey design was approved by the Princeton University IRB. All statistics published from those communications are in aggregate, with no personal identifiers attached to individual author responses. Datasets used for this work were already publicly available or were obtained with permission from study authors.
8 Open science
Code and data used for our analyses, as well as our full paper corpus, are included in our online SI (available at https://anonymous.4open.science/r/sok_misinformation-41E8).